mathslover:

Tutorial Drill: #1 Linear Inequalities

mathslover:

LINEAR INEQUATIONS : I have seen many people getting confused in this topic : Linear inequations : Inequations : A statement involving variable(s) and the sign of inequality viz. > , < ,≥ ,≤ is called an inequality or an inequation An equation can be linear, cubic , qudratic etc. and it may contain 1 or more than 1 variables. for example: \(\large{3x-2<0}\) \(\large{2x+3} \ge 0\) Solutions of an inequation : A solution of an inequation is the value (s)of the variable (s) that makes it a true statement . \(\large{\frac{3-2x}{5}=\frac{x}{3}-4}\) Let us take the above equation as an example : LHS of this inequation is \(\cfrac{3-2x}{5}\) and RHS is \(\cfrac{x}{3}-4\) we observe that: For x = 9 , We have LHS = \(\cfrac{3-2*9}{5} = -3\) and RHS : \(\cfrac{9}{3} - 4 = -1\) Clearly : -3<-1 LHS < RHS which is true. Hence x = 9 is one of the solution of this inequation. Similarly we can verify that no. greater than 7 is a solution of the given inequation. Solving a linear inequation : Properties of inequalities Let a, b and c be real numbers. Transitive Property If a < b and b < c then a < c Addition Property If a < b then a + c < b + c Subtraction Property If a < b then a - c < b - c Multiplication Property If a < b and c is positive then c*a < c*b If a < b and c is negative c*a > c*b Note: If each inequality sign is reversed in the above properties, we obtain similar properties. If the inequality sign < is replaced by <= ( less than or equal) or the sign > is replaced by >= ( greater than or equal ), we also obtain similar properties. Examplex : 1 1) 2x - 4 \(\le\) 0 2x - 4 + 4 \(\le\) 4 2x \(\le\) 4 x \(\le\) 2 [Properties of Linear Inequalities Info. has been taken from internet]

mathslover:

Tutorial Drill #1 is a part of my (mathslover's) tutorials that he contributed on OpenStudy. To make those available to all, I'm trying to repost them here, for the user's benefit. Thanks!

mathslover:

Link to this tutorial on OpenStudy: http://openstudy.com/study#/updates/4ff513f9e4b01c7be8c83b1d

Ultrilliam:

Would you like me to re-open this question for you?

Ultrilliam:

(So you can have both tutorials open at once :))

mathslover:

Yeah, if you can! :) It would be great.

Ultrilliam:

Gladly ^_^

Ultrilliam:

Done :)

mathslover:

Thanks man!

MARC:

Interesting. *thumbs up* \(:) \)

mathslover:

Thanks, @MARC ! :)

MARC:

^_^

pooja195:

Well done :)

KrissyBoy:

THANK. YOU. HOLY. LORD.

200082741:

TYYYYYYYYYYY

VAN1LLA:

This helps a lot !